# Efficient Generative Model Training via Embedded Representation Warmup

### v1

@@ 4 Embedded Representation Warmup: A Two-Phase Training Framework
Diffusion models generate high-fidelity images [16, 8], yet their training is impeded by the simultaneous challenges of learning robust semantic features and performing generative denoising [49]. This coupling creates an inherently complex optimization landscape that demands prolonged training periods.
In this section, we present Embedded Representation Warmup (ERW), a framework that strategically decouples these challenges into two distinct phases. In Phase 1, we initialize the early layers of the diffusion model with high-quality semantic features from pretrained models; in Phase 2, we transition to standard diffusion training with a gradually diminishing representation alignment term, allowing the model to increasingly focus on generation. This separation addresses the inefficiency of joint optimization in conventional diffusion models and significantly enhances training efficiency.
@@ 4.1 Phase 1: Representation Warmup Stage
Let $\mathcal{H}_{\mathbf{\theta}}(\mathbf{x})$ denote an encoder that maps an image $\mathbf{x}\in\mathcal{X}$ to its latent representation $\mathbf{z}\in\mathcal{Z}$, and let $\mathbf{f}_{\mathrm{rep}}:\mathcal{X}\to\mathcal{R}$ be a high-quality pretrained representation model. Our objective is to learn a mapping $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}$ such that its output closely aligns with $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$.
Specifically, we define the alignment loss as:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\Bigl{[}\bigl{\|}\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\bigr{\|}_{2}^{2}\Bigr{]}$ (13)
By minimizing Eq. (13), we ensure that $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}$ produces features well-aligned with the pretrained representation model. This alignment mitigates the need for the diffusion model to learn complex semantic features purely from noisy inputs; instead, subsequent layers can operate on features that are already semantically meaningful.
@@ 4.2 Phase 2: Generative Training with Decaying Representation Guidance
In this phase, the alignment loss $\mathcal{L}_{\mathrm{align}}$ encourages the diffusion model’s hidden representations, after being projected by a trainable function $\mathcal{T}_{\mathbf{\theta}}$, to maintain alignment with the high-quality semantic features of the external encoder:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{t,\mathbf{x}}\Biggl{[}\Bigl{\|}\mathcal{T}_{\mathbf{\theta}}\Bigl{(}\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))\Bigr{)}-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\Bigr{\|}_{2}^{2}\Biggr{]}$ (15)
@@
4.3 Theoretical Analysis: Warmup Leads to Faster Convergence
Our warmup phase minimizes the misalignment between the model’s internal representation $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))$ and the target semantic features $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$. After warmup, this misalignment becomes negligible for almost all data samples $\mathbf{x}$.
With the residual error minimized, the overall training objective during the diffusion stage is effectively dominated by the generative loss $\mathcal{L}_{\mathrm{diffusion}}$. The alignment loss $\mathcal{L}_{\mathrm{align}}$, scaled by the diminishing weight $\lambda(t)$, contributes minimally to the total loss. This enables the model to focus on refining its generative performance rather than correcting representation discrepancies.
To formalize this intuition, consider the gradient of the total loss with respect to the model parameters $\mathbf{\theta}$:
$\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{total}}=\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{diffusion}}+\lambda(t)\cdot\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{align}}$ (17)
Since the alignment loss has been minimized during warmup, its gradient $\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{align}}$ is small. Moreover, the alignment weight $\lambda(t)$ decays over time (as specified in Eq. (16)), further reducing the impact of the second term. Consequently, parameter updates are predominantly driven by the diffusion loss:
$\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{total}}\approx\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{diffusion}}$ (18)
This separation means the model no longer needs to allocate capacity to learn semantic features from scratch during the main training phase. Instead, it can concentrate on perfecting its generative capabilities, leading to substantially faster convergence and improved sample quality.
@@ 5 Experiments
5.1 Setup
$\bullet$ Implementation Details. We adhere closely to the experimental setups described in DiT [40] and SiT [36], unless otherwise noted. Specifically, we utilize the ImageNet dataset [7], preprocessing each image to a resolution of $256\times 256$ pixels. Following the protocols of ADM [8], each image is encoded into a compressed latent vector $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE [44]. For our model configurations, we employ the B/2, L/2, and XL/2 architectures as introduced in the SiT papers, which process inputs with a patch size of 2. To ensure a fair comparison with SiT models, we maintain a consistent batch size of 256 throughout training. Further experimental details, including hyperparameter settings and computational resources, are provided in Appendix C.
$\bullet$ Evaluation. We report Fréchet inception distance (FID; Heusel et al. 15), sFID [37], inception score (IS; Salimans et al. 46), precision (Pre.) and recall (Rec.) [27] using 50K samples. We also include CKNNA [19] as discussed in ablation studies. Detailed setups for evaluation metrics are provided in Appendix D.
Table 1: System-level comparison on ImageNet 256$\times$256 with CFG. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. Results that include additional CFG scheduling are marked with an asterisk (*), where the guidance interval from [28] is applied for ERW.
@@ 5.1 Setup
  $0$+ ERW (ours) 0040 01.94 7.91 300.9 0.75 0.63
$\bullet$ Sampler and Alignment objective. Following SiT [36], we always use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the number of function evaluations (NFE) as 250 by default. We use Normalized Temperature-scaled Cross Entropy (NT-Xent) training objective for alignment.
$\bullet$ Baselines. We use several recent diffusion-based generation methods as baselines, each employing different inputs and network architectures. Specifically, we consider the following four types of approaches: (a) Pixel diffusion: ADM [8], VDM+++++ + [23], Simple diffusion [18], CDM [17], (b) Latent diffusion with U-Net: LDM [44], (c) Latent diffusion with transformer+U-Net hybrid models: U-ViT-H/2 [2], DiffiT [12], and MDTv2-XL/2 [11], and (d) Latent diffusion with transformers: MaskDiT [55], SD-DiT [56], DiT [40], and SiT [36]. Here, we refer to Transformer+U-Net hybrid models that contain skip connections, which are not originally used in pure transformer architecture. Detailed descriptions of each baseline method are provided in   Appendix E.
5.2 Comparison to Prior Methods
Table 2: FID comparisons with SiT-XL/2. In this table, we report the FID of ERW with SiT-XL/2 on ImageNet $256\times 256$ at various Training iterations. Here is only full training without warmup, because we load a well trained warmuped checkpoint. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar iterations or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without classifier-free guidance.
@@ 5.2 Comparison to Prior Methods
+ERW (ours) 675M 100K 6.0 6.37 207.5 0.67 0.63
Figure 4: Comparison of Training Efficiency and Cost Analysis with Warmup and Full Training Stages. Left: Scatter plot depicting the relationship between the total training cost (in TFLOPs) and the FID score for various training strategies. Each point is annotated with a simplified label (e.g., “10K+90K”) representing the warmup and full training split, and the marker sizes are scaled based on a combination of the FID and FLOPs values to highlight the relative differences. Right: Bar chart comparing the computational costs of the warmup and full training stages for different strategies (all evaluated over 100K iterations). The chart shows the warmup cost, full training cost, and their corresponding total cost.
Table 1 summarizes our results on ImageNet $256\times 256$ under classifier-free guidance (CFG). Our ERW significantly boosts the convergence speed of SiT-XL/2, enabling strong FID scores at just 40 epochs. Importantly, we observe that ERW provides an alternative path to large-scale diffusion modeling by separating representational learning from generative training, thereby reducing the need for extended epochs.   Figure 3 illustrates generated samples on ImageNet, further confirming the high-quality outputs achieved by ERW.
@@ 5.3 Effect of ERW: Quantitative Gains and Training Efficiency
We begin by examining how ERW influences SiT-XL/2’s Frechet Inception Distance (FID) when no classifier-free guidance (CFG) is used.
@@
    Efficient FID Improvements. In  Table 5.2, ERW consistently achieves competitive or superior FID values compared to baselines. For instance, ERW reaches an FID of 6.0 at 100k iterations, markedly outperforming the REPA method [54] within the same iteration budget.
@@
Next, we analyze how splitting the total training budget between warmup and full diffusion training impacts both generation quality and computational overhead ( Figure 4).
  1. (a)
    Training Splits. We experiment with various allocations: 10k warmup + 90k full training, 20k warmup + 80k full training, and so on, always keeping the total at 100k iterations.
  2. (b)
    Accelerated Convergence. As shown in  Figure 4, a 10k warmup already yields a notable FID improvement (down to 6.4) with only a slight increase in total FLOPs ( Figure 4). This underscores that early alignment with a strong pretrained encoder can significantly hasten the later diffusion training.
Table 3: Analysis of ERW on ImageNet 256$\times$256. All models are SiT-B/2 trained for 50K iterations. All metrics except FID without classifier-free guidance. For Acc., we report linear probing results on the ImageNet validation set using the latent features aligned with the target representation. We fix $\lambda=0.5$ here. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively.
@@
Table 4: Analysis of ERW places influences in SiT-B/2. All models are based on SiT-B/2 and trained for 50K iterations under the batch size of 256 without using classifier-free guidance on ImageNet $256\times 256$. $\downarrow$ indicates lower values are better. Results empirically validate our hypothesis that placing ERW at the forefront of the architecture yields optimal performance.
@@
Table 5: Analysis of ERW depth, projection depth, and different dynamic or consistent projection loss $\lambda$ influences in SiT-XL/2. All models are based on SiT-XL/2 and trained for 50K iterations under a batch size of 256 without using classifier-free guidance on ImageNet $256\times 256$. The target representation model is DINOv2-B, and the objective is NT-Xent. $\downarrow$ indicates lower values are better. The results show that a projection depth of 14 and a projection loss $\lambda$ of 4.0 yield substantial improvements in both FID and sFID, indicating an optimal configuration for model performance.
[…]
SiT-XL/2 + REPA [54] 52.3 31.24 24.3 0.45 0.53
3 8 0.5 13.3 13.65 113.2 0.69 0.48
4 8 0.5 11.7 17.93 114.5 0.68 0.48
5 8 0.5 13.1 15.25 117.0 0.69 0.48
6 8 0.5 13.7 16.62 115.7 0.68 0.49
8 8 0.5 15.6 16.42 104.0 0.67 0.46
12 8 0.5 19.2 16.91 87.8 0.65 0.48
4 8 0.5 13.7 17.93 114.5 0.68 0.48
4 10 0.5 12.8 17.25 121.2 0.68 0.48
4 12 0.5 12.3 17.03 123.7 0.69 0.48
4 14 0.5 12.4 16.97 124.4 0.68 0.48
4 16 0.5 12.6 16.79 124.2 0.67 0.49
4 14 0.1 21.6 19.58 74.9 0.62 0.49
4 14 0.5 12.4 16.97 124.3 0.68 0.48
4 14 1.0 10.5 15.07 139.9 0.70 0.48
4 14 2.0 09.0 11.17 151.2 0.72 0.51
4 14 4.0 08.6 09.84 153.6 0.72 0.51
4 14 6.0 08.8 09.46 151.1 0.71 0.52
5.4 Ablation Studies
Target Representation.
@@ 5.4 Ablation Studies
We hypothesize that early layers in the diffusion backbone primarily learn semantic features (the L2R circuit), whereas deeper layers specialize in generative decoding.
@@
    Empirical Validation. In  Table 5.3, initializing the earliest layers (0–8) notably outperforms re-initializing middle or late sections (FID 54.2 vs. $>67$).
@@
We also investigate how the final projection head depth and the alignment-loss coefficient $\lambda$ affect training ( Table 5.3).
@@
    Optimal Hyperparameters. Using 4 warmup layers, a projection head at depth 14, and $\lambda=4.0$ achieves an FID of 8.6 at 50k iterations—a substantial gain over baselines.
@@
Finally, we measure layer-wise representation quality using Class-conditional $k$-Nearest Neighbor Accuracy (CKNNA) [3], which indicates how well the hidden features capture class discriminability.
@@
(a) Alignment with ERW
(b) Training Dynamic for Alignment
Figure 5: Scalability of ERW. (a) Alignment with ERW plot showing that the representation model’s alignment with the ERW model exhibits a high degree of alignment across all ERW depths. (b) Training dynamics for alignment indicate that within the 500K training steps for SiT-XL/2, the alignment between DINOv2-g and the diffusion model first decreases and then increases.

### v2

@@ 4 Embedded Representation Warmup: A Two-Phase Training Framework
Diffusion models generate high-fidelity images (Ho et al., 2020; Dhariwal & Nichol, 2021), yet their training is impeded by the simultaneous challenges of learning robust semantic features and performing generative denoising (Song et al., 2020). This coupling creates an inherently complex optimization landscape that demands prolonged training periods.
In this section, we present Embedded Representation Warmup (ERW), a framework that strategically decouples these challenges into two distinct phases. In Phase 1, we initialize the early layers of the diffusion model with high-quality semantic features from pretrained models; in Phase 2, we transition to main diffusion training with a gradually diminishing representation alignment term, allowing the model to increasingly focus on generation. This separation addresses the inefficiency of joint optimization in conventional diffusion models and significantly enhances training efficiency.
@@ 4.1 Phase 1: Representation Warmup Stage
Let $\mathcal{H}_{\mathbf{\bm{\theta}}}(\mathbf{x})$ denote an encoder that maps an image $\mathbf{x}\in\mathcal{X}$ to its latent representation $\mathbf{z}\in\mathcal{Z}$, and let $\mathbf{f}_{\mathrm{rep}}:\mathcal{X}\to\mathcal{R}$ be a high-quality pretrained representation model. Our objective is to learn a mapping $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}$ such that its output closely aligns with $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$.
Specifically, we define the warmup loss as:
$\mathcal{L}_{\mathrm{warmup}}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\Bigl{[}\bigl{\|}\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\bm{\theta}}}(\mathbf{x}))-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\bigr{\|}_{2}^{2}\Bigr{]}$ (13)
By minimizing Eq. (13), we ensure that the early layers of the diffusion model, denoted $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}$, produce features on clean latents that are well-aligned with the pretrained representation model. This alignment mitigates the need for the diffusion model to learn complex semantic features purely from noisy inputs; instead, subsequent layers can operate on features that are already semantically meaningful.
@@ 4.2 Phase 2: Generative Training with Decaying Representation Guidance
In this phase, the alignment loss $\mathcal{L}_{\mathrm{align}}$ encourages the diffusion model’s hidden representations for noisy inputs to align with the external encoder’s features for clean inputs, following REPA. A trainable projection head $\mathcal{T}_{\mathbf{\bm{\theta}}}$ maps the intermediate features to the representation space:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{t,\mathbf{x},\mathbf{\epsilon}}\Biggl{[}\Bigl{\|}\mathcal{T}_{\mathbf{\bm{\theta}}}\Bigl{(}\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\Bigr{)}-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\Bigr{\|}_{2}^{2}\Biggr{]}$ (15)
where $\mathcal{R}_{\mathbf{\bm{\theta}}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)$ is the intermediate representation from the early layers of the model given noisy latent $\mathbf{z}_{t}$.
@@ 5 Experiments
Table 1: System-level comparison on ImageNet 256$\times$256 with CFG. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. Results marked with an asterisk (*) use advanced CFG scheduling techniques; specifically, for our method, we apply the guidance interval scheduling from (Kynkäänniemi et al., 2024).
@@
  000+ REPA* 0800 01.42 4.70 305.7 0.80 0.65
  $0$+ ERW (ours) 0200 01.64 4.71 260.2 0.78 0.66
  $0$+ ERW (ours)* 0350 01.41 4.46 293.9 0.79 0.65
5.1 Setup
$\bullet$ Implementation Details. We adhere closely to the experimental setups described in DiT (Peebles & Xie, 2023) and SiT (Ma et al., 2024), unless otherwise noted. Specifically, we utilize the ImageNet dataset (Deng et al., 2009), preprocessing each image to a resolution of $\times 256$ 256 pixels. Following the protocols of ADM (Dhariwal & Nichol, 2021), each image is encoded into a compressed latent vector $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE (Rombach et al., 2022). For our model configurations, we employ the B/2 and XL/2 architectures as introduced in the SiT papers, which process inputs with a patch size of 2. To ensure a fair comparison with SiT models and REPA, we maintain a consistent batch size of 256 throughout training. Further experimental details, including hyperparameter settings and computational resources, are provided in Appendix C.
$\bullet$ Evaluation. We report Fréchet inception distance (FID; Heusel et al. 2017), sFID (Nash et al., 2021), inception score (IS; Salimans et al. 2016), precision (Pre.) and recall (Rec.) (Kynkäänniemi et al., 2019) using 50K samples. We also include CKNNA (Huh et al., 2024) as discussed in ablation studies. Detailed setups for evaluation metrics are provided in Appendix D.
$\bullet$ Sampler and Alignment objective. Following SiT (Ma et al., 2024), we always use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the number of function evaluations (NFE) as 250 by default. We use Normalized Temperature-scaled Cross Entropy (NT-Xent) training objective for alignment.
Table 2: FID comparisons with SiT-XL/2. In this table, we report the FID of ERW with SiT-XL/2 on ImageNet $\times 256$ 256 at various Training iterations. This table shows results from the full training phase, starting from a checkpoint pre-trained with our warmup phase. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar iterations or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without Classifier-Free Guidance.
@@ 5.1 Setup
+ERW (ours) 675M 100K 12.1 5.25 94.2 0.69 0.63
$\bullet$ Baselines. We use several recent diffusion-based generation methods as baselines, each employing different inputs and network architectures. Specifically, we consider the following four types of approaches: (a) Pixel diffusion: ADM (Dhariwal & Nichol, 2021), VDM+++++ + (Kingma & Gao, 2024), Simple diffusion (Hoogeboom et al., 2023), CDM (Ho et al., 2022), (b) Latent diffusion with U-Net: LDM (Rombach et al., 2022), (c) Latent diffusion with transformer+U-Net hybrid models: U-ViT-H/2 (Bao et al., 2023), DiffiT (Hatamizadeh et al., 2024), and MDTv2-XL/2 (Gao et al., 2023), and (d) Latent diffusion with transformers: MaskDiT (Zheng et al., 2024), SD-DiT (Zhu et al., 2024), DiT (Peebles & Xie, 2023), and SiT (Ma et al., 2024). Here, we refer to Transformer+U-Net hybrid models that contain skip connections, which are not originally used in pure transformer architecture. Detailed descriptions of each baseline method are provided in   Appendix E.
5.2 Comparison to Prior Methods
Figure 4: Comparison of Training Efficiency and Cost Analysis with Warmup and Full Training Stages. Bar chart comparing the computational costs of the warmup and full training stages for different strategies. The chart shows the warmup cost, full training cost, and their corresponding total cost.
Table 1 summarizes our results on ImageNet $\times 256$ 256 under Classifier-Free Guidance (CFG). Our ERW significantly boosts the convergence speed of SiT-XL/2, enabling strong FID scores at 350 epochs.   Figure 3 illustrates generated samples on ImageNet, further confirming the high-quality outputs achieved by ERW.
@@ 5.3 Effect of ERW: Quantitative Gains and Training Efficiency
We begin by examining how ERW influences SiT-XL/2’s FID when no CFG is used.
@@
    Efficient FID Improvements. In  Table 5.1, ERW consistently achieves competitive or superior FID values compared to baselines. For instance, ERW reaches an FID of 12.1 with 100k warmup + 100k full training, markedly outperforming the REPA method (Yu et al., 2024) within the same iteration budget.
@@
Next, we analyze how splitting the total training budget between warmup and full diffusion training impacts both generation quality and computational overhead. As shown in  Figure 4, FLOPs for warmup is fewer than the full training.
Table 3: Analysis of ERW on ImageNet 256$\times$256. All models are SiT-B/2 trained for 50K iterations. All metrics except FID without Classifier-Free Guidance. We fix $\lambda=0.5$ here. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively.
@@
5.4 Ablation Studies
Target representation.
Table 4: Analysis of ERW places influences in SiT-B/2. All models are based on SiT-B/2 and trained for 50K iterations under the batch size of 256 without using Classifier-Free Guidance on ImageNet $\times 256$ 256. $\downarrow$ indicates lower values are better. Results empirically validate our hypothesis that placing ERW at the forefront of the architecture yields optimal performance.
@@ 5.4 Ablation Studies
Table 5: Analysis of ERW depth, projection depth, and different dynamic or consistent projection loss $\lambda$ influences in SiT-XL/2. All models are based on SiT-XL/2 and trained for 100K iterations under a batch size of 256 without using Classifier-Free Guidance on ImageNet $\times 256$ 256. The target representation model is DINOv2-B, and the objective is NT-Xent. $\downarrow$ indicates lower values are better. The results show that a projection depth of 14 and a projection loss $\lambda$ of 4.0 yield substantial improvements in both FID and sFID, indicating an optimal configuration for model performance.
[…]
SiT-XL/2 + REPA  (Yu et al., 2024) 19.4 6.06 67.4 0.64 0.61
3 8 0.5 14.4 5.28 82.7 0.68 0.62
4 8 0.5 13.8 5.31 87.1 0.68 0.62
5 8 0.5 13.4 5.29 87.8 0.68 0.63
6 8 0.5 13.6 5.30 87.3 0.67 0.63
8 8 0.5 15.4 5.37 82.3 0.66 0.63
12 8 0.5 16.2 5.36 79.2 0.66 0.63
5 10 0.5 12.9 5.29 90.4 0.68 0.63
5 12 0.5 12.5 5.24 92.0 0.69 0.62
5 14 0.5 12.5 5.26 91.5 0.69 0.63
5 16 0.5 12.3 5.25 93.4 0.69 0.62
5 18 0.5 12.1 5.25 94.2 0.69 0.63
5 20 0.5 12.6 5.27 92.3 0.69 0.63
5 18 0.1 16.6 5.31 75.8 0.67 0.60
5 18 1.0 12.7 5.41 92.8 0.68 0.64
5 18 2.0 13.3 5.39 90.5 0.68 0.63
5 18 4.0 13.1 5.38 92.2 0.68 0.64
5 18 6.0 13.4 5.45 91.6 0.67 0.63
@@
We hypothesize that early layers in the diffusion backbone primarily learn semantic features (the L2R circuit), whereas deeper layers specialize in generative decoding. The placement of the alignment loss is therefore critical. We specify the alignment target using "Depth X-Y", which means the alignment loss is computed on the output of layer Y, using a projection head that takes features from layers X through Y as input.
@@
    Empirical Validation. In  Table 5.4, initializing the earliest layers (0–8) notably outperforms re-initializing middle or late sections (FID 54.2 vs. >67>67> 67).
@@
We also investigate how the final projection head depth and the alignment-loss coefficient $\lambda$ affect training ( Table 5.4). The projection head, $\mathcal{T}_{\mathbf{\bm{\theta}}}$, is a deep MLP that, similar to REPA, maps intermediate transformer features to the dimensionality of the target representation.
@@
    Optimal Hyperparameters. Using 5 warmup layers, a projection head at depth 18, and $\lambda=0.5$ achieves an FID of 12.1 at 100k iterations—a substantial gain over baselines.
@@
Figure 5: Scalability of ERW. Training dynamics for alignment indicate that within the 500K training steps for SiT-XL/2, the alignment between DINOv2-g and the diffusion model first decreases and then increases.
Finally, we measure layer-wise representation quality using Class-conditional $k$-Nearest Neighbor Accuracy (CKNNA) (Caron et al., 2021), which indicates how well the hidden features capture class discriminability.